2 runners can average 8 and 10 yards per second respectively. if the slower runner was given 7 1/2 second head start in a race that ended in a tie, what was the distance they raced?
Answer by mananth(16949)

runner A = 8 yards/second
Runner B = 10 yards/second
..
Runner A 7.5 seconds head start
When B starts he has run 7.5 *8 =60 yards.
..
the make up speed between them is 10-8 = 2 yards /second
..
distance = 60 yards
..
t= 60/2
t = 30 second.
..
Runner B catches up runner A after 30 seconds at the finishing line.
rate of runner B = 10 yards/second
..
Distance = 10*30 = 300 yards.
..
Race distance = 300 yards
..
m.ananth@hotmail.ca